OpenStudy (anonymous):

A physics student throws a ball with initial speed V0 = 10 m/s at 45 degrees. The ball strikes the wall 3.0 m away a. When will it strike the wall? Answer = .43s b. Is the ball ascending/descending? c. What is the height at which the ball strikes the wall? (With respect to the initial height) Answer: 2m d. Find the impact velocity of the ball Answer: 7i+2.9j

OpenStudy (anonymous):

I think his answer is wrong for a possibly. \[Vy = 10\sin45, Vx = 10\cos45\] So I used the equation:\[Y = Vyi(t)−1/2(g)(t^2)\] \[3 = 10\sin45(t) -1/2(10)(t^2) = -5t^2+7.1t-3\] but I got t = .34 s

OpenStudy (anonymous):

Im not sure about d., does anyone know how he got 7i+2.9j?

OpenStudy (badhi):

They've given horizontal displacement as 3, not vertical. Use \(S=ut\) horizontally instead. For d, horizontal component of the velocity will not change. Find the vertical velocity component at the impact by applying appropriate equations vertically.

OpenStudy (anonymous):

a) you need to consider horizontal movement: \[v_x=v_0·\cos(\theta) \rightarrow s_x=v_0·\cos(\theta)·t \rightarrow t_{impact}=\frac{ 3 }{ 10·\cos(45)}=0.42 \sec\]b) Peak height should happen when: \[v_y=v_0·\sin(\theta)-g t=0\rightarrow t_{PEAK}=\frac{ v_0·\sin(\theta) }{ g }=0.72\sec\]That means that before reaching peak height, it has hit the wall, therefore it was ascending c)Height is given by:\[h(t)=v_0·\sin(\theta)·t-\frac{ 1 }{ 2 }g t^2\rightarrow h(t=0.42)=2.12m\] d)Velocity is given by: \[\bar v(t)=v_0·\cos(\theta) \hat i+[v_0\sin(\theta)-g·t] \hat j\] just replace t with the impact time (0.42 sec) and get the vector

OpenStudy (anonymous):

Thanks! that makes sense